OpenStudy (anonymous):

sec^2x+3=tan^2x+4 verify the identities

OpenStudy (anonymous):

Note the following identity: \(\bf sec^2(x)=tan^2(x)+1\).

OpenStudy (anonymous):

Now we can change either side of the equation to verify the identity.

OpenStudy (anonymous):

???^

OpenStudy (anonymous):

OK first we must choose which side we will change so that it looks like the other side of the equation. I will choose the right hand side (one can choose left side if they wish). Now replace the \(\bf tan^2(x)\) with \(\bf sec^2(x)-1\) to get:\[\bf R.H.S=\tan^2(x)+4=(\sec^2(x)-1)+4=\sec^2(x)+3=L.H.S\]And so the right and left hand sides of the equations are equal. @Gabysolis49 understand?